Recent Log Entries

Follow-up Re: South Linac u-tubes

Log Number: 3854102Logbook(s): Body: 

Done, except that primary supply and return utubes were not put back in 2L20 or 2L21. This will wait until SRF is ready to cool the modules back down.

iocsfeldec restoration failed

Log Number: 3854101Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was The output from the myRestore execution was as follows: <MyAPI>: Channel lookup fault [iocsfeldec:ioc_sr.VALD]

injector penetration drop distance

Log Number: 3854100Logbook(s): Entry Makers: gramesBody: 

Pete asked what is the cable drop distance from ISB to injector.

Courtesy of John Hansknecht, Fig. 1 shows it's a 34 foot drop distance from ISB floor to tunnel floor, and of course the shorter distance of 25 feet to the elevated injector cable tray.

Note the pre-injector tunnel floor is about 18"? lower than the NL.

iocitf2 restored

Log Number: 3854097Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-10-05 14:06:52 The output from the myRestore execution was as follows: iocitf2 caen6533@2020-10-05 14:06 108(108) restored blm2@2020-10-05 14:06 48(48) restored systemEnum@2020-10-05 14:06 2(2) restored RasterUITF@2020-10-05 14:06 3(3) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ued -A -s -i iocitf2 (success) Procedure return status: Success

South Linac u-tubes

Log Number: 3854092Logbook(s): Entry Makers: brianmBody: 

We are working on removing L-tubes from south linac return transfer line at 2L03,14,20,21 and installing u-tubes in 2L14,20,21.

Raster 4000 - center slice

Log Number: 3854088Logbook(s): Entry Makers: sandorfiBody: 

- continuing raster 4000 study,
- slice through center of spot w lights - fig 1
- slice through center of spot -no lights - fig 2
- horizontal slice w raster OFF - fig 3

check raster 4000

Log Number: 3854087Logbook(s): Entry Makers: sandorfiBody: 

IBC-transfer @20 A;
IBC-Main @ 59 A;
Dump sol @ 320 A;

- "sweet" raster using AWF with 4000 pts (from last week)
- Period = 1.5 ms; elliptical coeff = 1.09
- Vpp(H) =0.1603; Vpp(V) =0.1475
- I(H) = 1.058 A; I(V) = 0.883
=> M905 ( fig 1)
=> dumpA, lights on (fig 2)
=> dumpA, lights off (fig 3)
ref: raster OFF (fig 4)

SLinac LCW cleanup Update

Log Number: 3854086Logbook(s): Body: 

INST and FML did a great job getting the LCW off the floor of the SLinac.

It is all cleaned up.

centering in 900 series quads, easy to steer through IBC now that copper target is removed

Log Number: 3854079Logbook(s): Entry Makers: poelkerBody: 

With the copper target removed, it's not hard at all to center the beam on viewer ITVM905 and put beam on the dump viewer in the desired location. No "dogleg" required. Will we regret centering the beam on ITVM905, or should we come into IBC low and left? Must ponder.

BUrt compare shows differences today, beam through IBC with and without the copper target.

Quads were adjusted to provide beam on viewers that is more circular. Had to adjust the 804 dipole to quad center, correctors were adjusted accordingly. Then note 904 and 904A related to ITV905 viewer.

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Log Number: 3854077Logbook(s): Body: 

vented beamline from gun valve to chopper.
Disconnected at each girder, blanked off for transport
Removed gun 2 and 3 laser windows for evaluation.
Blanked y chamber and roughed out, pump on local supply.
Powered UHV ion pump supplies for 1I and 0I regions off and disconnected cables.
Disconnected suitcase, roughed out tee and ion pump on local supply.
Removed PCup cross for easier handling.

All instrumentation, air and water disconnected from y chamber through chopper.

three days to perform target work and restore beam

Log Number: 3854076Logbook(s): Entry Makers: poelkerBody: 

On Friday, Neil Ricky and Bern removed the rooftiles, then Team HDIce rotated the cryostat vertical and removed the copper alignment target, then rotated the cryostat horizontal and restored vacuum, then Neil Ricky and Bern replaced the rooftiles, got the RadCon seal of approval, and beam restored to dump by Monday noon. Not bad, and excellent practice for when we install/replace targets during October, November and December.

injector UHV supplies

Log Number: 3854075Logbook(s): Entry Makers: marcyBody: 

powered off UHV supplies from 1I01 through chopper, disconnected at patch panel. Girders appear to be disconnected and free for removal.

Pages